Select the correct Null and Alternative Hypothesis for the following situation: A consumer group claims that 590 out of 1,000 insurance claims are settled within 30 days. An insurance company states that this is not the proportion of claims that are settled within 30 days.

a) Ho: p ≠ .5900 Ha: p > .5900

b) Ho: p = .5900 Ha: p ≠ .5900

c) Ho: p ≠ .5900 Ha: p = .5900

d) Ho: p ≠ .5900 Ha: p < .5900
Answer #1

P = X / n = 590 / 1000 = 0.59

H0 :- P = 0.59

H1 :- P 0.59

Option b is correct.